OpenStudy (anonymous):

A solid has 14 faces and 12 vertices. How many edges does the solid have? A. 18 edges B. 20 edges C. 22 edges D. 24 edges

OpenStudy (anonymous):

24

OpenStudy (texaschic101):

formula : number of faces + number of vertices minus the number of edges = 2. It can be written like this : F + V - E = 2 lets plug in our numbers... 14 + 12 - E = 2 26 - E = 2 26 - 2 = E 24 = E And that is how you do that :)

OpenStudy (anonymous):

thanks! I forgot about tht method.